Schwarz, a German firm, is considering a four- year project in China that requires an investment (capital expenditure) of EUR 600,000. The following additional information is presented:
R t = CNY 1,750,000 x 1.15 t -1 (i. e., annual growth rate of revenue is 15 percent)
Direct expenses are 25 percent of revenues.
Fixed expenses are CNY 250,000 annually.
Net working capital for each year equals 25 percent of revenues for the following year. Taxes are 32 percent.
Assume straight- line depreciation for tax purposes. After- tax salvage value is CNY 4 million. Spot EURCNY equals 7.0. Assume that interest rates in Europe and China are 1.5 percent and 6 percent, respectively, and annual compounding. Making appropriate assumptions, what is the EUR cash flows in year 2?
Answer #1

Solution :

Calculation of EUR Cash flows for year 2

Revenue for year 2 ( CNY 1,750,000 * 1.15 ) CNY 2,012,500

Less : Direct Expense ( 25% of 1,750,000 ) CNY 503,125

Less: Fixed Expense CNY 250,000

Less: Net Working Capital [ 25% of ( 1,750,000 * 1.15 * 1.15 ) ] CNY 578,593.75

Cash flows before depreciation and tax = CNY 680,781.25

Cash flows Before Depreciation and tax in EUR = CNY 680,781.25 / 7.31 = EUR 93,130.13

Less : Depreciation ( EUR 600,000 / 4 )     = EUR 150,000

Cash Flows Before Tax = ( EUR 56,869.87 )

Add: Tax Saved @ 32 % =   EUR 18,198.36

Cash flows After Tax = ( EUR 38,671.51 )

Add : Depreciation    = EUR 150,000

Cash flows for year 2 = EUR 111,328.49

Working Note :

Calculation of Forward Rate

Forward Rate ( CNY/EUR ) = 7 * ( 1 + 0.06 ) / ( 1 + 0.015 ) = CNY 7.31 per EUR